what Celsius temperature is required to increase the volume of a sealed balloon from 2.0 L to 3.0 L if the initial temperature is 15°C and the atmospheric pressure is constant?
Answer #1

Solution

We know that Charles law states that, at a constant pressure the volume of an ideal gas is proportional to the temperature.

V \alpha T

V = kT

V/T = k

So,

V1/T1 = V2/​​​​​​T2

Where, V1 = 2.0 L, V2 = 3.0 L, T1 = 15°C = 288. 15 K, T2 = ?

Putting these values in the above equation then we get

2.0 L/288.15 K = 3.0 L/T2

T2 = 288.15 × 3.0 L/ 2.0

T2 = 432.225 K = 159.075°C

159.075°C temperature is required to increase the volume of a sealed balloon from 2.0 L to 3.0 L.